A research team has suggested that men like Pokemon more then women. A group of 270 women and a group of 302 men were asked if they like Pokemon. Out of the group of 270 women, 120 like Pokemon while 216 men like Pokemon.

(a) Calculate the best point estimate for the true value by which the proportion of men that like Pokemon more then women.

(b) Use the data to perform a test of hypothesis at a significance level of 5% of the claim that the sex of the individual effects if they like 5 Pokemon. [Include the critical region, the value of the pivot for the sample data and the corresponding p-value]
Answer #1

a)

The difference in point estimate = 216/302 - 120/270 = 0.7152 - 0.4444 = 0.2708

b)

For sample 1, we have that the sample size is Ni = 270, the number of favorable cases is Xi 120, so then the sample proportio

(5) Conclusion It is concluded that the null hypothesis Ho is rejected. Therefore, there is enough evidence to claim that pop